OpenStudy (anonymous):

Help g(x) = 4 + x + e^x; Find inverse of g(5) or g^-1(5) I got: y = 4 + x + e^x x = 4 + y + e ^y x - 4 = y + e^y and...What do I do with the e^y? Thank you.

hartnn (hartnn):

finding g^-1 (5) means to put g(x)= 5 and finding the value of x. so, 5= 4 + x + e^x can u find one obvious solution to this ?

OpenStudy (anonymous):

really? is that how it is? I tohught you have to ge the inverse first then substitute 5

hartnn (hartnn):

thats the simpler way out..

OpenStudy (anonymous):

you could take ln to both sides. This will make "e" disappear and the exponent with be brought down as a coefficient. ln e = 1

OpenStudy (anonymous):

so.. x - 4 = y + e^y ln x - ln 4 = ln y + y log (base e) x - log (base e) 4 = log(base e) y + y.. ...how would i isolate y? andharnm...how would it work out if i followed your suit. would it still give me the same answer as the long method?

OpenStudy (anonymous):

hartnn*

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

u cannot isolate y by taking log, all u can do is guess one of the obvious solution as x=0 verify it.

OpenStudy (anonymous):

I couldn't think of a way either. hopefully someone else will come along

OpenStudy (anonymous):

...so.. 5 = 4 + x + e^x 5 = 4 + (0) + (e)^0 5 = 4 + 0 + 1 5 = 5 so...what does that tell me... that there is a solution somewhere along the way?

hartnn (hartnn):

this tells you that g^-1 (5) = 0 and since g is a function., g^-1 (5) has a unique solution and in this case that unique solution is 0

OpenStudy (anonymous):

ok..so the answe r is 0 then?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

hahahahaha!! you're right! Cool!!! Thanks.

hartnn (hartnn):

welcome :)